? ? A metal rod 1.5 m long rotates about its one end in a vertical plane at right angles to the magnetic meridian. If the frequency of rotation is 20 rev/s. Find the emf induced between the ends of rod. B H = 0.32 ´ 10 -4 T: Horizontal component of earth’s magnetic field.

??

A metal rod 1.5 m long rotates about its one end in a vertical plane at right angles to the magnetic meridian. If the frequency of rotation is 20 rev/s. Find the emf induced between the ends of rod.


                             BH = 0.32 ´ 10-4T: Horizontal component of earth’s magnetic field.


Grade:

1 Answers

Shane Macguire
14 years ago

Dear aakash,

The given information is as follows:

Length of rod     = 1.5 mt.

                             Frequency of rotation    = 20 rev/sec.

                             BH = 0.32 * 10-4 T

                             Angular velocity of rod = 20 * 2∏ rad/sec.

                             emf induced = B*dA/dt = 0.32*10-4 *20*∏*(1.5)2.

       Therefore horizontal component of the earth's magnetic feild is  =  4.5 * 10-3 V.

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

Akhilesh Shukla